OpenStudy (anonymous):

Solve the equation. Give a general formula for all the solutions. 2tan4(θ)=-2√3

OpenStudy (anonymous):

general formula?

OpenStudy (anonymous):

oh damn, was that \[2\tan(4x)=-2\sqrt{3}\]? i must be tired same idea though make \[\tan(x)=-\sqrt{3}\] so \(4x=\frac{2\pi}{3}\) and \(x=\frac{\pi}{6}\)

OpenStudy (anonymous):

you should get something like \[x=\frac{(3k-1)\pi}{12}\] when you are done finding a few of them and generalizing